OpenStudy (anonymous):

simplify and find restrictions. x+1/x^2+2x-8 - x/4x-8

OpenStudy (anonymous):

\[\frac{ x+1 }{ x^2+2x-8}-\frac{ x }{ 4x-8 }\]

OpenStudy (anonymous):

the problem is, =(x+1)/(x^2+2x-8)-x/(4x-8) =(x+1)/[(x+4)(x-2)]-x/2(x-4) =[2(x+1)(x-4)-x(x+4)(x-2)]/[(x+4)(x-2)(x+4)] =(2x^2-6x-8-x^3-2x^2+8x)/[(x+4)(x-4)(x-2)] =(-x^3+2x-8)/[(x+4)(x-4)(x-2)] the restriction is the denominator should not be zero, i.e. x should not be +4,-4,+2